OpenStudy (anonymous):

If 125cm (squared) is the volume of a cube, how do you find the side length?

OpenStudy (solomonzelman):

all sides of the cube are equal so (lets say that s is the side of the cube) \[s^3=Volume~~~~(of~~~the~~cube)\] In your problem, \[s^3=125~~~~~~~->~~~~~~~~~s^3=5^3~~~~~~~~->~~~~~~~s=5\]

OpenStudy (anonymous):

Thank you!